OpenStudy (anonymous):

For the equation, 4x - 2y + z = 0, write the subspace of sulutions in R(3) as the span of two vectors in R(3).

OpenStudy (helder_edwin):

when u solve the equation u get: \[ \large x=\frac{1}{2}y-\frac{1}{4}z. \]

OpenStudy (helder_edwin):

then u construct a "solution" vector: \[ \large \begin{pmatrix} x\\ y \\ z \end{pmatrix}=\begin{pmatrix} 1/2y-1/4z\\ y \\ z \end{pmatrix} \]

OpenStudy (helder_edwin):

now split the vector into two vectors, one for each parameter.

OpenStudy (anonymous):

so y(1/2, 1, 0) and z(-1/4, 0, 1).... which is just what i would do after reduced echelon in any other matrix, which is what I thought, but the solution in the book states this... \[span \left\{ (1, 0, -4)^{t}, (0, 1, 2)^{t}\right\}\] I dont understand how they got that.

OpenStudy (helder_edwin):

u have \[ \large y(1/2,1,0)^t+z(-1/4,0,1) \] replace \(y=2\alpha\) and \(z=-4\beta\) and see what u get.

OpenStudy (anonymous):

\[\alpha(1, 2, 0)^{t}, \beta(1, 0, -4)^{t}\]

OpenStudy (anonymous):

but thats not the same thing is it?

OpenStudy (helder_edwin):

your book must have a typo. the second vector u got is the same as the first in the solution.

OpenStudy (anonymous):

okay I will work the next one and see what I get... so is writing \[span \left\{ (1, 2, 0)^{t}, (1, 0, -4)^{t} \right\}\] the same thing as writing it \[x=y \left(\begin{matrix}1 \\ 2\\ 0\end{matrix}\right) + z \left(\begin{matrix}1 \\ 0\\-4\end{matrix}\right)\]

OpenStudy (anonymous):

and also would it of made any difference if I solved for y or z instead of x in the beginning, even though the vectors would be different it would still be the right answer wouldn't it? is there any particular reason you chose x to start with.

OpenStudy (helder_edwin):

it is not the same thing. actually the answer to the problem is \[ \large \text{span}\{(1,2,0)^t,(1,0,-4)^t\}. \] and what u wrote last is wrong. don't forget \(x\) is NOT

OpenStudy (helder_edwin):

a vector

OpenStudy (helder_edwin):

regarding which variable u solve in terms of the others. it really makes no difference. of course the vectors u obtain will be different.

OpenStudy (anonymous):

thank you.

OpenStudy (helder_edwin):

u r welcome